# Parc Científic UV, c/ Catedrático José Beltrán, 2, E-46980 Paterna (València), Spain

### v2

@@ Abstract
Keywords:
Majorana Neutrino Masses, Casas-Ibarra parametrisation, (Linear/Inverse) Seesaw Mechanism, (Generalised) Scotogenic Model, Zee Model
1 Introduction
@@
The origin of the light neutrino masses remains one of the most interesting questions in particle physics. Several possible mechanisms have been proposed, one of the most compelling explanations being that neutrinos are Majorana particles, as hinted by the Weinberg operator Weinberg:1979sa . One of the simplest high-energy completions are heavy right-handed neutrinos, such that light neutrino masses are generated by the (Type-I) seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Yanagida:1979gs ; Glashow:1979nm ; Gell-Mann:1979vob . Other possibilities, like radiative neutrino mass models, have also been extensively explored Ma:1998dn ; Cai:2017jrq ; Gargalionis:2020xvt .
@@
To study the phenomenology and do parameter scans of neutrino mass models, it is very useful to parametrise the Yukawa couplings that appear in the neutrino mass matrix in terms of light neutrino parameters [^1] It is worth noting the existence of models in which lepton mixing is connected to quark mixing Minakata_2004 ; Giarnetti:2024vgs ; Harada_2006 ; PICARIELLO_2007 ; Chauhan_2007 ; Barranco:2010we ; Zhang:2012pv ; Zhukovsky:2019eoy , without the need for Grand Unification. In such frameworks, the PMNS matrix is expressed in terms of the quark mixing parameters, so that the Yukawa couplings depend on both neutrino and quark observables, combined in specific ways depending on the model. (masses, mixings and phases) and other free parameters of the models. For this goal, several parametrisations have been proposed in the literature. In particular, this approach has been very successfully applied to the Type-I seesaw model using the Casas-Ibarra (CI) parametrisation Casas:2001sr , studied for special Yukawa textures in Ref. Ibarra:2003xp , simplified in the case of just two right-handed neutrinos in Ref. Ibarra:2003up , and generalised to include radiative corrections Pilaftsis:1991ug in Lopez-Pavon:2015cga (see also Refs. Broncano:2003fq ; Guo:2006qa ; Xing:2007zj ; Xing:2011ur ; Heeck:2012fw ; Xing:2024cyc for other studies). In Refs. Blennow:2011vn ; Donini:2012tt , other exact parametrisations for models with right-handed neutrinos, valid outside the seesaw limit, have been derived. For Type-II seesaw Magg:1980ut ; Mohapatra:1980yp ; Lazarides:1980nt ; Schechter:1980gr , obtaining the (symmetric) Yukawa matrix in terms of neutrino masses is trivial, while for the Type-I+II seesaw a parametrisation has been proposed in Ref. Akhmedov:2008tb . Furthermore, parametrisations similar to the CI one for Type-III and inverse seesaw can be found in Refs. Das:2012ze ; Das:2020uer , a modified CI parametrisation for the KNT model Krauss:2002px was given in Ref. Cepedello:2020lul , and a generalised CI parametrisation for an extension of the scotogenic model Tao96 ; Ma:2006km with several inert doublets was proposed in Ahriche:2022bpx .
@@
It is worth noting that, although in many works the seesaw parameter space is explored by adopting the standard CI parametrisation – where several degrees of freedom of the model are embedded in an orthogonal complex matrix – this is not the only possible approach. Alternative formulations have been proposed in the literature, aiming to relate more directly the parameters of the high-energy Lagrangian to low-energy observables. In particular, Refs. Davidson:2001zk ; Ellis:2002fe ; Ibarra:2005qi have shown that the parametrisation can be reformulated in terms of the hermitian combination $Y^{\dagger}Y$ of the Yukawa couplings, which provides a complementary perspective and can be particularly advantageous in the supersymmetric seesaw. Driven by the same motivation, another interesting parametrisation for the inverse seesaw was presented in Kriewald:2024rlg .
@@
In addition, other formulations for the Yukawa structure in less minimal models have been proposed. In Ref. Babu:2002uu , a parametrisation of the $n_{L}\times n_{L}$ antisymmetric Yukawa matrix of the singly-charged scalar present in the Zee-Babu model Babu:1988ki , which appears quadratically, has been obtained using the fact that the latter has a zero eigenvalue and it has been extensively used in phenomenological studies AristizabalSierra:2006gb ; Nebot:2007bc ; Ohlsson:2009vk ; Herrero-Garcia:2014hfa ; Schmidt:2014zoa . Further, Ref. Felkl:2021qdn also using this fact, obtained some restrictions on the antisymmetric Yukawa matrix of models in which it appears linearly, like in the Zee model Zee:1980ai ; Cheng:1980qt ; Wolfenstein:1980sy (see also Ref. Chen:2025thp ). Finally, in Refs. Cordero-Carrion:2018xre ; Cordero-Carrion:2019qtu , another general but more involved parametrisation for Majorana neutrino mass models has been proposed.
@@
The rest of the paper is organised as follows. In Sec. 2 we describe the general neutrino mass matrix and the generalised CI parametrisation. In Sec. 3 we follow the form of the neutrino mass matrix and study typical model realisations of it, both at tree level and at one loop: i) the seesaw and the Scotogenic Model; ii) the linear seesaw and the Generalised Scotogenic Model; and iii) the linear+inverse seesaw and the Extended Scotogenic Model. In Sec. 4 we discuss models with extra symmetries and the application of the generalised CI parametrisation to the Zee Model. In Sec. 5 we discuss an approach to parametrise the Yukawas using the complete neutrino mass matrix in models with right-handed neutrinos. In App. A we derive the standard CI in the seesaw limit using this approach. Our conclusions are summarised in Sec. 6.
2 General neutrino mass matrix and parametrisation
@@
$m_{\nu}=Y_{1}^{T}M_{11}Y_{1}+Y_{1}^{T}M_{12}Y_{2}+Y_{2}^{T}M_{21}Y_{1}+Y_{2}^{T}M_{22}Y_{2}\,.$ (3)
The latter form may be mapped onto neutrino mass–model parameters, which we exploit in Sec. 3. The sum of these contributions can be rewritten as
@@
with $\mathcal{M}=\mathcal{M}^{T}$ being complex symmetric. Therefore, the neutrino mass matrix $m_{\nu}$ can always be expressed in the form of Eq. (4), regardless of the underlying physics or the order at which the different mass contributions arise. Note that this clearly also holds for $N>2$, using appropriate block matrices. This has already been pointed out in App. F of Ref. Cordero-Carrion:2019qtu . [^2] Eq. 1 of Refs. Cordero-Carrion:2018xre ; Cordero-Carrion:2019qtu presents the general Majorana neutrino mass matrix in a form that, while algebraically equivalent to ours, is less transparent when several diagonal and off-diagonal terms are present. In such cases - and when $y_{1}$ and $y_{2}$ in that equation are subject to specific constraints - a direct use of their master parametrisation does not follow automatically.
To study the phenomenology, it is useful to express the Yukawa couplings in terms of light neutrino parameters and other free parameters. For Type-I seesaw, this may be done via the CI parametrisation Casas:2001sr , which can be used in the seesaw limit (see Sec. 3.1.1 below). However, in the generic case, before being able to apply the CI procedure Casas:2001sr , the extended mass matrix $\mathcal{M}$ has to be diagonalised via the Autonne-Takagi factorisation
@@
where $D_{\mathcal{M}}$ is a real diagonal matrix with non-negative entries and $V$ is a unitary matrix, $V^{\dagger}V=\mathbb{I}_{n_{1}+n_{2}}$. Then exactly the same derivation as in Ref. Casas:2001sr can be applied to $V\mathcal{Y}$. Also by means of the Autonne-Takagi factorisation, the $n_{L}\times n_{L}$ light neutrino mass matrix may be written as
@@
where $\mathcal{R}$ is an $(n_{1}+n_{2})\times n_{L}$ (semi)-orthogonal matrix, $\mathcal{R}^{T}\mathcal{R}=\mathbb{I}_{n_{L}}$, and we assume that there are no strictly vanishing eigenvalues in $D_{\mathcal{M}}$. Here the word generalised highlights that, by writing the Majorana neutrino mass matrix as in Eq. (4) and diagonalising also $\mathcal{M}$, the structure of the CI parametrisation of Ref. Casas:2001sr actually holds for any Majorana neutrino mass model. [^3] A generalised CI parametrisation was proposed in Ahriche:2022bpx , but in the context of one specific model, namely a scotogenic model with several inert doublets (rather than one as in Tao96 ; Ma:2006km ), and using an orthogonal matrix in one of the diagonalisation steps, instead of the unitary matrix appearing in the Autonne-Takagi factorisation of Eq. (6).
Note that this expression is valid for any field basis. Indeed, one can work in a basis where the charged lepton mass matrix is not diagonal, as discussed, e.g., in Refs. Ibarra:2003xp ; Ibarra:2003up . In this case, the matrix $U$ can be written as $U=U_{\text{PMNS}}\;U_{L}$, where $U_{\text{PMNS}}$ is the standard PMNS mixing matrix Pontecorvo:1957cp ; Pontecorvo:1957qd ; Maki:1962mu ; Pontecorvo:1967fh and $U_{L}$ accounts for the charged lepton mixing. [^4] The choice of mass ordering (normal or inverted) becomes relevant only at this point, via the parametrisation of the PMNS matrix, $U_{\text{PMNS}}$, and the order of the eigenvalues in $D_{m}$. It is also possible to operate in a basis where, e.g., the Majorana mass matrix of the right-handed neutrinos in the Type-I seesaw is not diagonal, as in Ref. Ibarra:2003up . However, we emphasise that in the GCI parametrisation proposed in Eq. (8), the matrix $V$ is the one that diagonalises the extended mass matrix $\mathcal{M}$ and it is not necessarily related to a change of basis, as in Ibarra:2003up .
@@
Any $n_{L}\times n_{L}$ Majorana neutrino mass matrix, which may have an arbitrary number of contributions, $m_{\nu}=\sum^{N}_{i,j=1}Y_{i}^{T}M_{ij}Y_{j}\,$, (9) with Yukawas $Y_{i}$ of dimensions $n_{i}\times n_{L}$, may be always written in the usual seesaw way, $\boxed{m_{\nu}={\mathcal{Y}}^{T}\,{\mathcal{M}}\,{\mathcal{Y}}}$ (10) with $\mathcal{Y}$ an $(n_{1}+\ldots+n_{N})\times n_{L}$ Yukawa matrix, and $\mathcal{M}$ an $(n_{1}+\ldots+n_{N})\times(n_{1}+\ldots+n_{N})$ complex symmetric matrix. Therefore, the seesaw expression is the most general form for any Majorana (neutrino) mass matrix. A generalised Casas-Ibarra (GCI) parametrisation may be used, $\mathcal{Y}=V^{\dagger}D^{-1/2}_{\mathcal{M}}\,\mathcal{R}\,D_{\sqrt{m}}\,U^{\dagger}\,\qquad\qquad\text{GCI\,\,parametrisation}\,$, (11) where $D_{\mathcal{M}}=V^{\ast}\mathcal{M}V^{\dagger}$ and $\mathcal{R}$ is an orthogonal matrix, $\mathcal{R}^{T}\mathcal{R}=\mathbb{I}_{n_{L}}$. This is a simple and well-suited parametrisation for numerical scans relevant to phenomenological analyses. (11)
A word of caution is in order. Note that, if the Yukawa matrices are not general, extra conditions on the matrix $\mathcal{R}$ may need to be imposed. For instance, this is the case of antisymmetric Yukawa matrices, $f=-f^{T}$, which appear for instance in the Zee (linear in the antisymmetric Yukawa) Zee:1985id and Zee-Babu models (quadratic in the antisymmetric Yukawa coupling) Babu:1988ki , see also Ref. Cheng:1980qt . For the Zee-Babu case, another parametrisation which uses the fact that an antisymmetric matrix has a zero eigenvalue is much more straightforward to use for phenomenological studies Babu:2002uu ; AristizabalSierra:2006gb ; Nebot:2007bc ; Herrero-Garcia:2014hfa ; Felkl:2021qdn , while for the Zee model Zee:1985id (see Refs. Herrero-Garcia:2017xdu ; Heeck:2023iqc ; Babu:2019mfe for detailed studies of the phenomenology), in Sec. 4 we derive a parametrisation of $\mathcal{R}$ that ensures the antisymmetric condition.
2.1 Specific form of the (semi-)orthogonal Casas-Ibarra $\mathcal{R}$ matrix
@@
3 Implementations of the mass matrix and the Generalised Casas-Ibarra parametrisation
In the following, we study in turn tree-level and loop-level implementations of diagonal, off-diagonal and general neutrino mass matrices. We consider specific models that are widely discussed in the literature and illustrate how to implement for them the GCI parametrisation. In all cases we start from a basis in which the $M_{ij}$ matrices defined above are diagonal, and we use the notation $\overline{M}_{i}$ for the eigenvalues of $\mathcal{M}$. Also note that, when working in the basis in which the charged lepton mass matrix is diagonal, the matrix $U$ appearing in all the expressions we give below coincides with the PMNS matrix, $U_{\text{PMNS}}$. For the latest values of the oscillation parameters, see Refs. Esteban:2024eli ; NuFIT60 .
3.1 Diagonal mass matrix
@@
3.1.1 Tree level: the Seesaw case
The most famous example of $M_{11}\neq 0$ at tree level is the usual seesaw limit of models with extra right-handed neutrinos. If $n_{R}=n_{1}$ right-handed neutrinos $N$ are added to the SM, the Yukawa and mass terms in the Lagrangian read Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Yanagida:1979gs ; Glashow:1979nm :

### v3

# Generalised Casas-Ibarra Parametrisation for Majorana Neutrino Masses

@@ Abstract
I Introduction
@@
The origin of the light neutrino masses remains one of the most interesting questions in particle physics. Several possible mechanisms have been proposed, one of the most compelling explanations being that neutrinos are Majorana particles, as hinted by the Weinberg operator [94]. One of the simplest high-energy completions are heavy right-handed neutrinos, such that light neutrino masses are generated by the (Type-I) seesaw mechanism [78, 100, 101, 50, 46]. Other possibilities, like radiative neutrino mass models, have also been extensively explored [71, 23, 44].
@@
To study the phenomenology and do parameter scans of neutrino mass models, it is very useful to parametrise the Yukawa couplings that appear in the neutrino mass matrix in terms of light neutrino parameters [^1] It is worth noting the existence of models in which lepton mixing is connected to quark mixing [77, 49, 54, 85, 28, 13, 104, 105], without the need for Grand Unification. In such frameworks, the PMNS matrix is expressed in terms of the quark mixing parameters, so that the Yukawa couplings depend on both neutrino and quark observables, combined in specific ways depending on the model. (masses, mixings and phases) and other free parameters of the models. For this goal, several parametrisations have been proposed in the literature. In particular, this approach has been very successfully applied to the Type-I seesaw model using the Casas-Ibarra (CI) parametrisation [24], studied for special Yukawa textures in Ref. [62], simplified in the case of just two right-handed neutrinos in Ref. [63], and generalised to include radiative corrections [86] in [70] (see also Refs. [20, 52, 97, 98, 56, 99] for other studies). In Refs. [16, 38], other exact parametrisations for models with right-handed neutrinos, valid outside the seesaw limit, have been derived. For Type-II seesaw [74, 81, 69, 91], obtaining the (symmetric) Yukawa matrix in terms of neutrino masses is trivial, while for the Type-I+II seesaw a parametrisation has been proposed in Ref. [6]. Furthermore, parametrisations similar to the CI one for Type-III and inverse seesaw can be found in Refs. [34, 33], a modified CI parametrisation for the Krauss-Nasri-Trodden (KNT) model [67] was given in Ref. [27], and a generalised CI parametrisation for an extension of the scotogenic model [93, 72] with several inert doublets was proposed in [3].
@@
It is worth noting that, although in many works the seesaw parameter space is explored by adopting the standard CI parametrisation – where several degrees of freedom of the model are embedded in an orthogonal complex matrix – this is not the only possible approach. Alternative formulations have been proposed in the literature, aiming to relate more directly the parameters of the high-energy Lagrangian to low-energy observables. In particular, Refs. [35, 39, 65] have shown that the parametrisation can be reformulated in terms of the hermitian combination $Y^{\dagger}Y$ of the Yukawa couplings, which provides a complementary perspective and can be particularly advantageous in the supersymmetric seesaw. Driven by the same motivation, another interesting parametrisation for the inverse seesaw was presented in [68].
@@
In addition, other formulations for the Yukawa structure in less minimal models have been proposed. In Ref. [10], a parametrisation of the $n_{L}\times n_{L}$ antisymmetric Yukawa matrix of the singly-charged scalar present in the Zee-Babu model [11], which appears quadratically, has been obtained using the fact that the latter has a zero eigenvalue and it has been extensively used in phenomenological studies [7, 82, 84, 59, 92]. Further, Ref. [42] also using this fact, obtained some restrictions on the antisymmetric Yukawa matrix of models in which it appears linearly, like in the Zee model [102, 30, 95] (see also Ref. [29]). Finally, in Refs. [31, 32], another general but more involved parametrisation for Majorana neutrino mass models has been proposed.
@@
The rest of the paper is organised as follows. In Sec. II we describe the general neutrino mass matrix and the generalised CI parametrisation. In Sec. III we follow the form of the neutrino mass matrix and study typical model realisations of it, both at tree level and at one loop: i) the seesaw and the Scotogenic Model; ii) the linear seesaw and the Generalised Scotogenic Model; and iii) the linear+inverse seesaw and the Extended Scotogenic Model. In Sec. IV we discuss models with extra symmetries and the application of the generalised CI parametrisation to the Zee Model. In Sec. V we discuss an approach to parametrise the Yukawas using the complete neutrino mass matrix in models with right-handed neutrinos. In App. A we derive the standard CI in the seesaw limit using this approach. Our conclusions are summarised in Sec. VI.
II General neutrino mass matrix and parametrisation
@@
$m_{\nu}$ $=Y_{1}^{T}M_{11}Y_{1}+Y_{1}^{T}M_{12}Y_{2}$
$+Y_{2}^{T}M_{21}Y_{1}+Y_{2}^{T}M_{22}Y_{2}\,.$ (3)
The latter form may be mapped onto neutrino mass–model parameters, which we exploit in Sec. III. The sum of these contributions can be rewritten as
@@
with $\mathcal{M}=\mathcal{M}^{T}$ being complex symmetric. Therefore, the neutrino mass matrix $m_{\nu}$ can always be expressed in the form of Eq. (4), regardless of the underlying physics or the order at which the different mass contributions arise. Note that this clearly also holds for $N>2$, using appropriate block matrices. This has already been pointed out in App. F of Ref. [32]. [^2] Eq. 1 of Refs. [31, 32] presents the general Majorana neutrino mass matrix in a form that, while algebraically equivalent to ours, is less transparent when several diagonal and off-diagonal terms are present. In such cases - and when $y_{1}$ and $y_{2}$ in that equation are subject to specific constraints - a direct use of their master parametrisation does not follow automatically.
To study the phenomenology, it is useful to express the Yukawa couplings in terms of light neutrino parameters and other free parameters. For Type-I seesaw, this may be done via the CI parametrisation [24], which can be used in the seesaw limit (see Sec. III.1.1 below). However, in the generic case, before being able to apply the CI procedure [24], the extended mass matrix $\mathcal{M}$ has to be diagonalised via the Autonne-Takagi factorisation
@@
where $D_{\mathcal{M}}$ is a real diagonal matrix with non-negative entries and $V$ is a unitary matrix, $V^{\dagger}V=\mathbb{I}_{n_{1}+n_{2}}$. Then exactly the same derivation as in Ref. [24] can be applied to $V\mathcal{Y}$. Also by means of the Autonne-Takagi factorisation, the $n_{L}\times n_{L}$ light neutrino mass matrix may be written as
@@
where $\mathcal{R}$ is an $(n_{1}+n_{2})\times n_{L}$ (semi)-orthogonal matrix, $\mathcal{R}^{T}\mathcal{R}=\mathbb{I}_{n_{L}}$, and we assume that there are no strictly vanishing eigenvalues in $D_{\mathcal{M}}$. Here the word generalised highlights that, by writing the Majorana neutrino mass matrix as in Eq. (4) and diagonalising also $\mathcal{M}$, the structure of the CI parametrisation of Ref. [24] actually holds for any Majorana neutrino mass model. [^3] A generalised CI parametrisation was proposed in [3], but in the context of one specific model, namely a scotogenic model with several inert doublets (rather than one as in [93, 72]), and using an orthogonal matrix in one of the diagonalisation steps, instead of the unitary matrix appearing in the Autonne-Takagi factorisation of Eq. (6).
Note that this expression is valid for any field basis. Indeed, one can work in a basis where the charged lepton mass matrix is not diagonal, as discussed, e.g., in Refs. [62, 63]. In this case, the matrix $U$ can be written as $U=U_{\text{PMNS}}\;U_{L}$, where $U_{\text{PMNS}}$ is the standard PMNS mixing matrix [88, 87, 75, 89] and $U_{L}$ accounts for the charged lepton mixing. [^4] The choice of mass ordering (normal or inverted) becomes relevant only at this point, via the parametrisation of the PMNS matrix, $U_{\text{PMNS}}$, and the order of the eigenvalues in $D_{m}$. It is also possible to operate in a basis where, e.g., the Majorana mass matrix of the right-handed neutrinos in the Type-I seesaw is not diagonal, as in Ref. [63]. However, we emphasise that in the GCI parametrisation proposed in Eq. (8), the matrix $V$ is the one that diagonalises the extended mass matrix $\mathcal{M}$ and it is not necessarily related to a change of basis, as in [63].
@@
Any $n_{L}\times n_{L}$ Majorana neutrino mass matrix, which may have an arbitrary number of contributions, $m_{\nu}=\sum^{N}_{i,j=1}Y_{i}^{T}M_{ij}Y_{j}\,$, (9) with Yukawas $Y_{i}$ of dimensions $n_{i}\times n_{L}$, may be always written in the usual seesaw way, $\boxed{m_{\nu}={\mathcal{Y}}^{T}\,{\mathcal{M}}\,{\mathcal{Y}}}$ (10) with $\mathcal{Y}$ an $(n_{1}+\ldots+n_{N})\times n_{L}$ Yukawa matrix, and $\mathcal{M}$ an $(n_{1}+\ldots+n_{N})\times(n_{1}+\ldots+n_{N})$ complex symmetric matrix. Therefore, the seesaw expression is the most general form for any Majorana (neutrino) mass matrix. A generalised Casas-Ibarra (GCI) parametrisation may be used, $\mathcal{Y}=V^{\dagger}D^{-1/2}_{\mathcal{M}}\,\mathcal{R}\,D_{\sqrt{m}}\,U^{\dagger}\,~\text{GCI\,\,parametrisation}\,$, (11) where $D_{\mathcal{M}}=V^{\ast}\mathcal{M}V^{\dagger}$ and $\mathcal{R}$ is a complex orthogonal matrix, $\mathcal{R}^{T}\mathcal{R}=\mathbb{I}_{n_{L}}$. This is a simple and well-suited parametrisation for numerical scans relevant to phenomenological analyses.
A word of caution is in order. Note that, if the Yukawa matrices are not general, extra conditions on the matrix $\mathcal{R}$ may need to be imposed. For instance, this is the case of antisymmetric Yukawa matrices, $f=-f^{T}$, which appear for instance in the Zee (linear in the antisymmetric Yukawa) [103] and Zee-Babu models (quadratic in the antisymmetric Yukawa coupling) [11], see also Ref. [30]. For the Zee-Babu case, another parametrisation which uses the fact that an antisymmetric matrix has a zero eigenvalue is much more straightforward to use for phenomenological studies [10, 7, 82, 59, 42], while for the Zee model [103] (see Refs. [60, 55, 9] for detailed studies of the phenomenology), in Sec. IV we derive a parametrisation of $\mathcal{R}$ that ensures the antisymmetric condition.
II.1 Specific form of the (semi-)orthogonal Casas-Ibarra $\mathcal{R}$ matrix
@@
III Implementations of the mass matrix and the Generalised Casas-Ibarra parametrisation
In the following, we study in turn tree-level and loop-level implementations of diagonal, off-diagonal and general neutrino mass matrices. We consider specific models that are widely discussed in the literature and illustrate how to implement for them the GCI parametrisation. In all cases we start from a basis in which the $M_{ij}$ matrices defined above are diagonal, and we use the notation $\overline{M}_{i}$ for the eigenvalues of $\mathcal{M}$. Also note that, when working in the basis in which the charged lepton mass matrix is diagonal, the matrix $U$ appearing in all the expressions we give below coincides with the PMNS matrix, $U_{\text{PMNS}}$. For the latest values of the oscillation parameters, see Refs. [41, 83].
III.1 Diagonal mass matrix
@@
III.1.1 Tree level: the Seesaw case
The most famous example of $M_{11}\neq 0$ at tree level is the usual seesaw limit of models with extra right-handed neutrinos. If $n_{R}=n_{1}$ right-handed neutrinos $N$ are added to the SM, the Yukawa and mass terms in the Lagrangian read [78, 100, 46, 101, 50]: